OpenStudy (anonymous):

question: A line segment AB has endpoints at A(5,9) and B(-7, -15). What is the sum of the coordinates of its midpoint?

OpenStudy (anonymous):

-4 i think

OpenStudy (anonymous):

HOw do you figure?

OpenStudy (anonymous):

first find the midpoint: 5-7/2=-1 9-15/2=-3 then add them : -1-3=-4

OpenStudy (anonymous):

our answer key says you are correct IA-just unsure how to get it!

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

why do you divide by 2?

OpenStudy (anonymous):

because you have to find the mid point, so the x and y values of the mid point are half of the distance between the two end points of the line

OpenStudy (anonymous):

thanks, that makes sense

OpenStudy (anonymous):